OpenStudy (anonymous):

Tangent to y = e^x log(2, x) at the point (1, 0)

OpenStudy (anonymous):

\[e^x(\log_{2}x) \]

zepdrix (zepdrix):

Mmm so what's the problem? Having trouble taking the derivative of a log that isn't base e?

OpenStudy (anonymous):

No, I get this as the tangent line equation \[\frac{ ex }{ \log2 }-\frac{ e }{ \log2 }\] but apparently my math program considers it incorrect

zepdrix (zepdrix):

Hmm ok let's see what's going on. \[\Large y=e^x\frac{\ln x}{\ln 2}\] \[\Large y'=\frac{1}{\ln2}\left(e^x \ln x+e^x\frac{1}{x}\right)\]

zepdrix (zepdrix):

Evaluating the function at x=1 gives us,\[\Large y'(1)=\frac{1}{\ln2}\left(\cancel{e^1\ln1}+e\right)\] Simplifies nicely to give us the slope of our tangent line,\[\Large m=\frac{e}{\ln2}\]

zepdrix (zepdrix):

\[\Large y_{\tan} \quad=\quad mx+b \qquad\to\qquad y_{\tan}\quad=\quad \frac{e}{\ln2}x+b\]

zepdrix (zepdrix):

Plugging in our point gives ussssssssss, Hmm ya it looks like you did everything correctly :\ strange...

OpenStudy (anonymous):

yeah, that's where I'm super confused.... Maybe you can help me out with another problem that I seem to be doing correctly as well, yet keep getting told I'm wrong: If $20,000 is invested in a savings account offering 3.5% per year, compounded continuously, how fast is the balance growing after 5 years? (Round your answer to the nearest cent.) I'm getting $764.99

OpenStudy (anonymous):

By the looks of it, it seems like a simple pert formula where the initial investment is subtracted and then the outcome is divided by 5 years to find the rate at that year, but apparently I am wrong?

zepdrix (zepdrix):

I think what we're doing is looking for the slope of the line tangent to the function A=Pe^rt at t=5. So we want \(\Large A'(5)\) What you described would be: The average rate of change across those 5 years. They're looking for the `instanteous rate of change` AT t=5 years. At least I think that's what they want, based on the wording.

OpenStudy (anonymous):

ok, I figured as much, but with having values for all the variable, show would I go about finding the derivative? Or do I find the derivative based on t and leave t empty so there is a variable?

zepdrix (zepdrix):

Correct, we want t `to vary` as we look for the derivative. We can't plug in t=5 and then take a derivative, that would mean we were holding t constant. Derivative, then plug 5. :)

OpenStudy (anonymous):

ok after derivatizing (coining this word) I get \[700e ^{.035t}\] Does that look correct?

zepdrix (zepdrix):

for A'? ya looks good so far! Now we're allowed to plug in 5.

OpenStudy (anonymous):

833.87 was right! Thanks for your help! Though we weren't able to understand why I was getting a no for the first question I'm glad you were able to help me out with my second!

zepdrix (zepdrix):

yay team \c:/ for the first one, perhaps they want you to round to a decimal for each value?

OpenStudy (anonymous):

I'll try that